# Generative model for financial time series trained with MMD using a signature kernel

### v3

@@ 5.6.3 Gain/loss Asymmetry and Leverage Effect
The negative skew in the observed returns can also be viewed through different lenses. The gain/loss asymmetry is the phenomenon that there are more large negative returns observed than large positive returns (see e.g.[12]). Figure 13(a) shows the expected ratio of positive returns (y-axis) given that the absolute value of the return exceeds a certain threshold (x-axis). For the threshold, we only use values up to 0.0237 which corresponds to the 95th percentile of the returns found in S&P 500 since there are limited samples beyond this point. On this measure, the MMD generated sequences are more aligned to the S&P 500 segments displaying a similar level of asymmetry. However, the COT-GAN generated sequences have a return distribution that is more symmetric. For the C-Sig-WGAN generated sequences, the asymmetry more or less follows the same pattern as the S&P 500 but is much more pronounced until it gets closer to the 95th percentile.
@@
The leverage effect is the observation that the volatility of the returns is negatively correlated with the returns (see e.g.[12]). We can quantify this by calculating the correlation between the returns and the squared returns in the subsequent days as shown in Figure 13(b). On this measure, we see that the MMD generated sequences only show a weak negative correlation in the immediate subsequent day before it quickly decays towards zero. This is in contrast to the S&P 500 segments and full sequence which shows a more pronounced negative correlation that is slower to decay. The COT-GAN generated sequences show a similar pattern to the MMD generated sequences while the C-Sig-WGAN exhibits a pronounced negative correlation in the immediate subsequent day before it quickly decays towards zero.
(a) Gain/loss asymmetry
(b) Leverage effect
(c) MMD vs COT-GAN for leverage effect
(d) MMD vs C-Sig-WGAN for leverage effect
@@
Figure 13: The top left plot shows the gain/loss asymmetry of the daily log returns. The y-axis is the expected ratio of positive returns given that the return exceeds a certain threshold on the x-axis. The other three plots show the correlation between the daily log returns and the squared returns of the subsequent day with lines and shaded regions being the same as in Figure 11.
@@ 6 Conclusion and Future Work
Several avenues exist for further exploration and improvement. The noise structure is clearly important for generating realistic sequences, and further research is needed to better understand the impact of different noise distributions on the generated data. The robustness of the generative model can be further improved by incorporating additional features or data sources to better capture the complexity of financial markets. For the RL application, one could introduce a Bayesian approach to learn the prevalent regime of the market and switch between different agents trained on different noise distributions or use an ensemble of weighted agents (see e.g. [14]). Expanding empirical studies and applications by handling multiple assets and investigating more sophisticated sampling and data preprocessing techniques are also possible directions for future research.
@@ Appendix B Moments of Returns
[…]

### v4

@@ 5.6.3 Gain/loss Asymmetry and Leverage Effect
The negative skew in the observed returns can also be viewed through different lenses. The gain/loss asymmetry is the phenomenon that there are more large negative returns observed than large positive returns (see e.g.[12]). Figure 13 shows the expected ratio of positive returns (y-axis) given that the absolute value of the return exceeds a certain threshold (x-axis). For the threshold, we only use values up to 0.0237 which corresponds to the 95th percentile of the returns found in S&P 500 since there are limited samples beyond this point. On this measure, the MMD generated sequences are more aligned to the S&P 500 segments displaying a similar level of asymmetry. However, both the NSDE and COT-GAN generated sequences have a return distribution that is more symmetric. For the C-Sig-WGAN generated sequences, the asymmetry more or less follows the same pattern as the S&P 500 but is much more pronounced until it gets closer to the 95th percentile.
@@
The leverage effect is the observation that the volatility of the returns is negatively correlated with the returns (see e.g.[12]). We can quantify this by calculating the correlation between the returns and the squared returns in the subsequent days as shown in Figure 14(a). On this measure, we see that the MMD generated sequences only show a weak negative correlation in the immediate subsequent day before it quickly decays towards zero. This is in contrast to the S&P 500 segments and full sequence which shows a more pronounced negative correlation that is slower to decay. The COT-GAN generated sequences show a similar pattern to the MMD generated sequences while the C-Sig-WGAN exhibits a pronounced negative correlation in the immediate subsequent day before it quickly decays towards zero. The NSDE generated sequences has the opposite effect with a positive correlation in the immediate subsequent day before it decays towards zero.
@@
Figure 13: The gain/loss asymmetry of the daily log returns. The y-axis is the expected ratio of positive returns given that the return exceeds a certain threshold on the x-axis.
(a) MMD
(b) NSDE
(c) COT-GAN
(d) C-Sig-WGAN
Figure 14: The plots show the correlation between the daily log returns and the squared returns of the subsequent $\tau$ days with lines and shaded regions being the same as in Figure 11.
5.6.4 Distance Based Comparison
Finally, we compare the generated sequences using two distance-based measures for time series distributions. First, we use $\widehat{\text{MMD}^{2}}$ to perform the permutation tests as in Section 4.3, but using the untruncated signature based on [46] for the signature kernel. This particular signature kernel is computed by numerically solving a related PDE and is computationally expensive. We choose a resolution for the grid such that the results do not change with further refinement. Second, we use the sliced adapted Wasserstein distance (SAW), also used in [1], which is computationally less expensive than the adapted Wasserstein distance.
There are 114 S&P 500 segments available and we extract corresponding sequences with the same time period to calculate both distances. Table 7 shows the results of the distance-based comparison. The MMD generated sequences perform the best for both distance measures, with the C-Sig-WGAN generated sequences coming second on both measures. Both the NSDE and COT-GAN generated sequences perform poorly on both measures.
Model $\widehat{\text{MMD}^{2}}$ p-value SAW mean SAW std
MMD 0.479 0.11 0.133 0.010
NSDE 3.561 0.0 0.153 0.010
COT-GAN 8.967 0.0 0.285 0.020
C-Sig-WGAN 0.762 0.05 0.137 0.007
Table 7: Distance based comparison of the generated sequences with the S&P 500 segments using $\widehat{\text{MMD}^{2}}$ based permutation test with a an untruncated signature kernel and sliced adpated Wasserstein distance (SAW).
@@ 6 Conclusion and Future Work
Several avenues exist for further exploration and improvement. The noise structure is clearly important for generating realistic sequences, and further research is needed to better understand the impact of different noise distributions on the generated data. The robustness of the generative model can be further improved by incorporating additional features or data sources to better capture the complexity of financial markets. For the RL application, one could introduce a Bayesian approach to learn the prevalent regime of the market and switch between different agents trained on different noise distributions or use an ensemble of weighted agents (see e.g. [14]). In our experiments, we found that RL tends to find an averaged strategy when mixing datasets from very different market environments. It requires additional architectural adjustments to produce distinct strategies for different environments (see [38]). Further research is needed to see how such adjustments can be generalised. Expanding empirical studies and applications by handling multiple assets and investigating more sophisticated sampling and data preprocessing techniques are also possible directions for future research.
Acknowledgments
J. Sester is grateful for financial support by the NUS Start-Up Grant Tackling model uncertainty in Finance with machine learning.
@@ Appendix B Moments of Returns
    The mean is annualised by business day convention which we calculate it as: $\textbf{Ann. returns}=252\times\bar{r}$.